# Maximizing Relation Extraction Potential: A Data-Centric Study to Unveil Challenges and Opportunities

### v1

@@ III-B2 Joint Relation Extraction Algorithms
[t] Task Type Algorithm Arch.1 Params.1 Super.1 Few.1 RC Recurrent Att-BLSTM BLSTM 2M $\checkmark$ PAWARE BLSTM 15M $\checkmark$ EntityAtt BLSTM 8M $\checkmark$ PLM R-BERT BERT 110M $\checkmark$ $\checkmark$ Roberta_ base RoBERTa (large) 355M $\checkmark$ $\checkmark$ LUKE RoBERTa (large) 483M $\checkmark$ ERNIE BERT enc+dec2 114M $\checkmark$ Prompt-PLM Know-Prompt RoBERTa (large) 355M $\checkmark$ $\checkmark$ GenPT RoBERTa (large) 355M $\checkmark$ Prompt-LLM GPTRE gpt-3.5-turbo NA3 $\checkmark$ Unleash-LLM gpt-3.5-turbo NA3 $\checkmark$ JRE SPN4RE BERT enc+dec2 110M $\checkmark$ TDEER BERT encoder 110M $\checkmark$ RIFRE BERT enc+dec2 109M $\checkmark$ UniRel BERT encoder 109M $\checkmark$
@@ III-C Experimental Methodology
This performance analysis aims to answer where, how, and why neural models fail to extract relations from textual data. For this, 15 neural relation extraction algorithms were selected to extract relations from the above-mentioned datasets. The algorithms were re-implemented using open-source codes available on GitHub. Algorithms such as PAWARE, Roberta_base, and UnleashLLM required entity type information, which was only available for TACRED and RETACRED datasets. Thus, the NER tool from Stanford CoreNLP [^14] https://stanfordnlp.github.io/CoreNLP/ [31] was used to annotate the other datasets with the entity type information. However, even this endeavor could not categorize all entities with their corresponding types, and many entities were annotated with the unknown (’O’) entity type. The statistics of this entity annotation process are available in Table XII in the Appendix. Also, the annotated datasets have been released with the source code. Finally, adjustments were made to the algorithms to accommodate for the incomplete entity-type information. Detailed implementation details of the affected algorithms can be found in Section B in the Appendix. The subsequent methodology can be divided into two categories: 1) Supervised and 2) Few-shot learning strategies.
@@ III-D Results
The trained algorithms were made to predict relationships from the test samples of the selected datasets. Tables XV, XVI and XVII in the Appendix depict the average performance and standard deviation for each algorithm dataset pair across five cross-validated folds (supervised relation classification and joint relation extraction) and three seed iterations (few-shot learning). This section highlights some findings from analyzing the performance of the algorithms to show the validation of the conducted experiments and highlight any intersections or deviations from trends in the literature. To this end, a critical evaluation of the relation classification and joint extraction algorithms has been discussed below.
@@ III-D1 Algorithms under fully-supervised setting
First, it can be inferred from Table XV that both PLM-based and Prompt-PLM algorithms were better at extracting relationships in the supervised setting than the recurrent algorithms. This finding matches the general trend in the literature where PLMs have shown significant performance gains at accurately extracting relationships from text. A probable reason for this observation can be due to the ability of the PLMs to understand the contextual information present in a sentence and come up with a comprehensive representation. However, it is worth mentioning that the performance of the recurrent architectures was not significantly behind that of the PLMs. This finding highlights the performance gap in the field, where even with highly advanced PLMs, not much performance gain was observed in the overall task of relation classification.
Next, the performances of the joint relation extractors were investigated based on the scores depicted in Table XVII. It was found that all joint relation extractors were extremely brittle and did not consistently perform across the datasets selected for this study. This finding can be observed by the high variation of performances, with most extractors performing the best on the WebNLG dataset and worst on the TACRED dataset. The existing literature on joint RE has only evaluated the algorithms on a few datasets. The observations from the experiments conducted in this study showcase that a significant performance gap exists in this domain, and it is essential to make joint extractors more generalizable.
@@ III-D2 Algorithms under low-resource setting
Next, in low resource settings, it was found that the Prompt-PLM algorithms consistently performed well on the relation classification task across the k-shots used for training/demonstrations as depicted by Table XVI. Two major inferences can be drawn from this finding. First, traditional PLM-based algorithms are inefficient at extracting relationships from textual data in low-resource scenarios, as seen by the low comparative performances of these algorithms. Second, converting the relation extraction objective to a language generation objective has significant benefits that best come into play without considerable training data.
Finally, the performances of the LLM-based algorithms were found to be highly volatile and depended on the properties of the target datasets. Although the algorithms DeepKE and GPT-RE showed performance gains for some dataset combinations, especially in the 1-shot setting, making these algorithms compatible with the task was difficult. First, the performance scores depicted in Table XVI for the LLM-based algorithms had to be calculated from a smaller subset of the test data because of the tendency of the LLMs to generate out-of-domain predictions that were not a part of the original label space. Thus, such samples had to be discarded from the performance calculation. This issue is evident by the extremely low performance on the TACRED and RETACRED datasets, which have complicated label space verbalization. Second, it was impossible to fit the full label space into the input prompt, especially for the larger dataset. Thus, the DeepKE algorithm could not be evaluated in a greater than 1-shot setting. This observation highlights the inefficiency of the LLM-based algorithms for relation classification, where it is common to have complex and large label spaces.
Despite having access to SOTA methods that caused performance to grow by leaps and bounds in several NLP areas, relation extraction, and classification are still far behind. Being the counterpart to algorithms, this naturally elucidates the necessity for a data-centric performance assessment to find possible attributes in the data that may hinder the performance of the relation extractors. The following section pursues this direction.
@@ IV A Data-centric Performance Comparison
This work investigates the attributes in the data that adversely affect the relation extraction capabilities of neural algorithms. To facilitate this objective, two types of analysis were conducted on the selected algorithm and datasets discussed in the previous section. First, a quantitative analysis was performed to critically evaluate the performance of the selected algorithms under various complex data characteristics. Second, a qualitative error analysis was conducted to reach the root cause of the incorrect predictions observed in the algorithms. To this end, the following sections discuss the findings from these analyses.
@@ IV-A Quantitative Analysis
This analysis was conducted to find connections between model performances and complex data characteristics. For this, the performance of the algorithms was observed for the complex and counter-complex categories discussed in Section III. Thus, the algorithms were analyzed under coarse vs. fine-grained relation distribution, short vs. long input, one-to-one vs. multiple-relations/overlap, and uniform vs. long-tail categories. Figures 4, 5, 7 and 8 show aggregate plots composed by averaging the performance scores of different algorithm-dataset combinations over five cross-validated folds (supervised learning) and three seed iterations (few-shot learning). Similarly, Figures 6 and 9 depict the average performances of joint relation extraction algorithms across five cross-validated training runs. The following sections discuss the findings from evaluating both relation classification and joint relation extractors with respect to the complex attributes.
@@
Similarly, distinct findings were observed with few-shot learning. First, it can be inferred from Figure 5(a) that Prompt-LLM algorithms showed superiority in tackling coarse relation distribution compared to PLM and Prompt-PLM algorithms. This finding further corroborates that the latter algorithms require a large sample space to learn relation distinctions accurately. On the other hand, the performance gain of the LLMs can be attributed to the vast amount of knowledge already present within these models, making them efficient at classification when the label space is sparse. However, the lower overall performance of the Prompt-LLMs in the fine-grained category suggests that LLMs tend to get confused when diverse relations are possible. Additionally, PLM and Prompt-PLM algorithms performed better at tackling fine-grained relations due to the aforementioned reasons. However, the large standard deviations experienced by these algorithms highlight that in low-resource settings, these algorithms are not robust at tackling fine-grained relations.
@@
Figure 4: Micro F1-score distribution of supervised relation classification
@@
Figure 5: Micro F1-score distribution of few-shot relation classification
@@
A fine-grained analysis was conducted to investigate the performance of the models by isolating different categories of multiple relations and overlapping entities. For the analysis of multiple relations, the data was divided into samples with only one relation ($n=1$), two relations ($n=2$), three relations ($n\geq 3$), four relations ($n\geq 4$), and five or more relations ($n\geq 5$). Similarly, for overlapping entities, the split included samples with no overlap (Normal), entity pair overlap (epo), and single entity overlap (seo), respectively. Tables IX, XI and X in the Appendix show the split statistics of the test data for datasets that suffer from this problem under supervised, few-shot and joint relation extraction domains.
@@
Finally, the complex characteristics of multiple relations and overlapping entities were investigated with respect to the performance of the joint relation extractors. It can be observed from Figure 9 that the performance of all extractors deteriorated as the number of relations increased in the given sample, with significant performance drops when more than 5 relations were present. Similarly, the performance of the relation extractors was significantly worse when tackling epo case. These two observations are interconnected as more relations would positively correlate with both entities sharing different relationships. The presence of multiple relations and overlapping entities should be tackled by the prediction of multiple triplets by the joint algorithms. However, the observations indicate that the algorithms were inefficient at doing so. A possible reason could be the discrepancy between the predicted and the ground truth triplets, which do not match under the ‘exact match’ setting.
@@
To facilitate the experiments, normalized weights were calculated for each relation label using min-max normalization with respect to the relational class consisting of the maximum and minimum samples. Figures 14(a)-14(c) in the Appendix show the heat maps of the F1 scores achieved by the relation classification algorithms for the relations of some of the long-tail datasets. It can be inferred from the figures that all relation classification and joint relation extraction algorithms were inefficient at handling relations with low representation in the datasets. This trend is evident from the dark-hued top rows corresponding to high F1 scores and the light-hued lowest rows correlating to low F1 scores on the heat maps. The problem stems from the inability of the models to learn an accurate representation of classes that do not have enough samples during training. Thus, these findings re-instantiate the prevalence of the long-tail distribution problem in the domain of neural relation extraction.
@@ IV-B Qualitative Error Analysis
All relation classification and joint relation extractors were employed in the supervised setting for this analysis. On the other hand, only the 1-shot experiments were used for few-shot learning, as the DeepKE algorithm was only compatible with 1-shot experiments. The test data and corresponding predictions from the models were combined. The RETACRED dataset was not used in this analysis, as it has the same text samples as TACRED. The final test samples were then segregated into simple and challenging classes according to the ease or difficulty of prediction. A voting-based scheme was devised to achieve this segregation. Therefore, the pipeline created by Alt et al. [21] based on erudite [77] was followed, and each data sample was scored based on the number of models correctly classifying it. To this end, the input samples were segregated as follows:
@@
The performance analysis highlighted the superiority of the PLM-based models when used in a supervised setting for relation classification. However, under certain conditions, the PLMs’ tendency to memorize the training distribution diffused their focus from the context of the input sample. Thus, this section investigated the context of the text samples to find possible reasons for this behavior. On manually inspecting the samples in the simple and challenging categories, it was found that most samples in the challenging class were difficult to understand, even for a human annotator. Two scenarios were found responsible for this “ambiguity”: 1) when the context provides insufficient or contradictory information, thereby confusing the model, and 2) when the relation definition is too fine-grained, and the model confuses between multiple relation labels. Some examples of the two cases of ambiguity are presented in Table V. To this end, the discussion presented below further corroborates these findings.
@@ V-A Ambiguity
The foremost challenge to supervised relation extraction is the aspect of contextual ambiguity in the input data. This issue was apparent from the large percentage of ambiguous samples in the challenging class for supervised and joint algorithms. The presence of ambiguous samples with diverging contextual information confuses the models, as the possibility of more than one correct relation exists. Furthermore, it was observed that relations with similar/correlating relation definitions posed a significant challenge for supervised relation extractors. Much like the previous case, these complicated relations led to the possibility of more than one correct relation per sample, and both supervised and joint extractors had difficulty handling them. Thus, alleviating these ambiguous scenarios would greatly help to make relation extractors more efficient. A possible future direction could be using low-resource algorithms such as prompt and LLM-based relation classifiers. As the previous sections demonstrated, these algorithms were not susceptible to ambiguity in context and relation labels. However, a more fine-grained analysis of the few-shot paradigm is needed, incorporating more algorithms and LLMs to get better insights into the workings of the algorithms. Additionally, adding external relation-specific information can be a possible way to tackle cases where the context does not provide sufficient information. Solving this complex scenario can have many benefits, as it can improve neural networks at fine-grained language understanding.
@@ V-B Joint Relation Extraction
Existing research in this field focuses on testing joint algorithms on a few datasets with similar characteristics. This practice has led to a narrow view of these algorithms, where they perform exceptionally well on the said datasets. According to the findings presented in this paper, one of the significant issues with joint extractors proved to be the false negative problem. This problem stems from the insufficiency of ground-truth labels, so even semantically correct predictions are deemed misclassifications. Thus, methods for creating better datasets with comprehensive annotations that do not require any manual labor are a need of the hour. Another strategy would be introducing enhanced evaluation schemes that do not entirely rely on ground-truth labels. Additionally, it was found that research on joint relation extraction using LLMs has been scarce in the literature. Thus, conducting exhaustive research in this direction is crucial to making joint relation extractors robust and generalizable.

### v2

@@ III-B2 Joint Relation Extraction Algorithms
[t] Task Type Algorithm Arch.1 Params. Super. Few. RC Recurrent Att-BLSTM BLSTM 2M $\checkmark$ PAWARE BLSTM 15M $\checkmark$ EntityAtt BLSTM 8M $\checkmark$ PLM R-BERT BERT 110M $\checkmark$ $\checkmark$ Roberta_ base RoBERTa (large) 355M $\checkmark$ $\checkmark$ LUKE RoBERTa (large) 483M $\checkmark$ ERNIE BERT enc+dec2 114M $\checkmark$ Prompt-PLM Know-Prompt RoBERTa (large) 355M $\checkmark$ $\checkmark$ GenPT RoBERTa (large) 355M $\checkmark$ Prompt-LLM GPTRE gpt-3.5-turbo NA3 $\checkmark$ Unleash-LLM gpt-3.5-turbo NA3 $\checkmark$ JRE SPN4RE BERT enc+dec2 110M $\checkmark$ TDEER BERT encoder 110M $\checkmark$ RIFRE BERT enc+dec2 109M $\checkmark$ UniRel BERT encoder 109M $\checkmark$
@@ III-C Experimental Methodology
This performance analysis aims to answer where, how, and why neural models fail to extract relations from textual data. For this, 15 neural relation extraction algorithms were selected to extract relations from the above-mentioned datasets. The algorithms were re-implemented using open-source codes available on GitHub [^7] https://github.com/. Algorithms such as PAWARE, Roberta_base, and UnleashLLM required entity type information, which was only available for TACRED and RETACRED datasets. Thus, the NER tool from Stanford CoreNLP [^8] https://stanfordnlp.github.io/CoreNLP/ [31] was used to annotate the other datasets with the entity type information. However, even this endeavor could not categorize all entities with their corresponding types, and many entities were annotated with the unknown (‘O’) entity type. The statistics of this entity annotation process are available in Table XII in the Appendix. Also, the annotated datasets have been released with the source code. Finally, adjustments were made to the algorithms to accommodate for the incomplete entity-type information. Detailed implementation details of the affected algorithms can be found in Section A-B in the Appendix. The subsequent methodology can be divided into two categories: 1) Supervised and 2) Few-shot learning strategies.
@@ III-D Results
The trained algorithms were made to predict relationships from the test samples of the selected datasets. Figures 4, 5 and 6 depict the average performance and standard deviation for each algorithm dataset pair across five cross-validated folds (supervised relation classification and joint relation extraction) and three seed iterations (few-shot learning). Detailed performance scores for the figures can be found in Tables XV, XVII and XVI in the Appendix. This section highlights some findings from analyzing the performance of the algorithms to show the validation of the conducted experiments and highlight any intersections or deviations from trends in the literature. To this end, a critical evaluation of the relation classification and joint extraction algorithms has been discussed below.
@@ III-D1 Algorithms under fully-supervised setting
First, it can be inferred from Figure 4 that both PLM-based and Prompt-PLM algorithms were better at extracting relationships in the supervised setting than the recurrent algorithms. This finding matches the general trend in the literature where PLMs have shown significant performance gains at accurately extracting relationships from text. A probable reason for this observation can be due to the ability of the PLMs to understand the contextual information present in a sentence and come up with a comprehensive representation. However, it is worth mentioning that the performance of the recurrent architectures was not significantly behind that of the PLMs. This finding highlights the performance gap in the field, where even with highly advanced PLMs, not much performance gain was observed in the overall task of relation classification.
Next, the joint relation extractors were investigated based on the performances depicted in Figure 5. It was found that all joint relation extractors were extremely brittle and did not consistently perform across the datasets selected for this study. This finding can be observed by the high variation of performances, with most extractors performing the best on the WebNLG dataset and worst on the TACRED dataset. The existing literature on joint RE has only evaluated the algorithms on a few datasets. The observations from the experiments conducted in this study showcase that a significant performance gap exists in this domain, and it is essential to make joint extractors more generalizable.
Figure 4: Average micro-F1 scores for all datasets and relation classifier combinations in the fully supervised setting. The x-axis has been magnified according to the range of scores achieved.
Figure 5: Average micro-F1 scores for all datasets and joint relation extractor combinations in the fully supervised setting.
@@ III-D2 Algorithms under low-resource setting
Next, in low resource settings, it can be inferred from Figure 6 that the Prompt-PLM algorithms consistently performed well across the k-shots in the traditional relation classification category. Two major inferences can be drawn from this finding. First, traditional PLM-based algorithms like RBERT and Roberta_base are inefficient at extracting relationships from textual data in low-resource scenarios, as seen by the low comparative performances of these algorithms. Second, converting the relation extraction objective to a language generation objective has significant benefits that best come into play without considerable training data.
Finally, the performances of the LLM-based algorithms were found to be highly volatile and depended on the properties of the target datasets. Although the algorithms UnleashLLM and GPT-RE showed performance gains for some dataset combinations, especially in the 1-shot setting, making these algorithms compatible with the task was difficult. First, the performance scores depicted in Figure 6 for the LLM-based algorithms had to be calculated from a smaller subset of the test data because of the tendency of the LLMs to generate out-of-domain predictions that were not a part of the original label space. Such samples had to be discarded from the performance calculation. This issue is evident by the extremely low performance on the WebNLG dataset, which has a complicated and large label space. Second, it was impossible to fit the full label space into the input prompt, especially for the larger dataset. Thus, the UnleashLLM algorithm could not be evaluated in more than 1-shot setting. This observation highlights the inefficiency of the LLM-based algorithms for relation classification, where it is common to have complex and large label spaces.
Figure 6: Average micro-F1 scores for all datasets and relation classifier combinations in the few-shot setting. Note that the UnleashLLM algorithm was only compatible with the k=1 setup.
Based on the findings discussed above, it can be inferred that there exist niche issues that hamper the performance of most relation extraction algorithms, even in the presence of SOTA architectures. Being the counterpart to algorithms, this naturally elucidates the necessity for a data-centric performance assessment to find possible attributes in the data that may hinder the performance of the relation extractors. The following section pursues this direction.
@@ IV A Data-centric Performance Comparison
This work investigates the attributes in the data that adversely affect the relation extraction capabilities of neural algorithms. To facilitate this objective, two types of analysis were conducted on the selected algorithm and datasets discussed in the previous section. First, a quantitative analysis was performed to critically evaluate the performance of the selected algorithms under various complex data characteristics. Second, a qualitative error analysis was conducted to reach the root cause of the incorrect predictions observed in the algorithms.
@@ IV-A Quantitative Analysis
This analysis was conducted to find connections between model performances and complex data characteristics. For this, the performance of the algorithms was observed for the complex and counter-complex categories discussed in Section III. Thus, the algorithms were analyzed under coarse vs. fine-grained relation distribution, short vs. long input, one-to-one vs. multiple-relations/overlap, and uniform vs. long-tail categories. Figures 7, 8, 10 and 11 show aggregate plots composed by averaging the performance scores of different algorithm-dataset combinations over five cross-validated folds (supervised learning) and three seed iterations (few-shot learning). Similarly, Figures 9 and 12 depict the average performances of joint relation extraction algorithms across five cross-validated training runs. The following sections discuss the findings from evaluating both relation classification and joint relation extractors with respect to the complex attributes.
@@
Similarly, distinct findings were observed with few-shot learning. First, it can be inferred from Figure 8(a) that Prompt-LLM algorithms showed superiority in tackling coarse relation distribution compared to PLM and Prompt-PLM algorithms. This finding further corroborates that the latter algorithms require a large sample space to learn relation distinctions accurately. On the other hand, the performance gain of the LLMs can be attributed to the vast amount of knowledge already present within these models, making them efficient at classification when the label space is sparse. However, the lower overall performance of the Prompt-LLMs in the fine-grained category suggests that LLMs tend to get confused when diverse relations are possible. Additionally, PLM and Prompt-PLM algorithms performed better at tackling fine-grained relations due to the above-mentioned reasons. However, the large standard deviations experienced by these algorithms highlight that in low-resource settings, these algorithms are not robust at tackling fine-grained relations.
@@
A fine-grained analysis was conducted to investigate the performance of the models by isolating different categories of multiple relations and overlapping entities. For the analysis of multiple relations, the data was divided into samples with only one relation ($n=1$), two relations ($n=2$), three relations ($n\geq 3$), four relations ($n\geq 4$), and five or more relations ($n\geq 5$). Similarly, for overlapping entities, the split included samples with no overlap (norm), entity pair overlap (epo), and single entity overlap (seo), respectively. Tables IX, XI and X in the Appendix show the split statistics of the test data for datasets that suffer from this problem under supervised, few-shot and joint relation extraction domains.
@@
Finally, the complex characteristics of multiple relations and overlapping entities were investigated with respect to the performance of the joint relation extractors. It can be observed from Figure 12 that the performance of all extractors deteriorated as the number of relations increased in the given sample, with significant performance drops when more than 5 relations were present. Similarly, the performance of the relation extractors was significantly worse when tackling samples with epo overlap. These two observations are interconnected as more relations would positively correlate with both entities sharing different relationships. The presence of multiple relations and overlapping entities should be tackled by the prediction of multiple triplets by the joint algorithms. However, the observations indicate that the algorithms were inefficient at doing so. A possible reason could be the discrepancy between the predicted and the ground truth triplets, which do not match under the ‘exact match’ setting.
@@
To facilitate the experiments, normalized weights were calculated for each relation label using min-max normalization with respect to the relational class consisting of the maximum and minimum samples. Figures 17(a)-17(c) in the Appendix show the heat maps of the F1 scores achieved by the relation classification algorithms for the relations of some of the long-tail datasets. It can be inferred from the figures that all relation classification and joint relation extraction algorithms were inefficient at handling relations with low representation in the datasets. This trend is evident from the dark-hued top rows corresponding to high F1 scores and the light-hued lowest rows correlating to low F1 scores on the heat maps. The problem stems from the inability of the models to learn an accurate representation of classes that do not have enough samples during training. Thus, these findings re-instantiate the prevalence of the long-tail distribution problem in the domain of neural relation extraction.
@@ IV-B Qualitative Error Analysis
All relation classification and joint relation extractors were employed in the supervised setting for this analysis. On the other hand, only the 1-shot experiments were used for few-shot learning, as the UnleashLLM algorithm was only compatible with 1-shot experiments. The test data and corresponding predictions from the models were combined. The RETACRED dataset was not used in this analysis, as it has the same text samples as TACRED. The final test samples were then segregated into simple and challenging classes according to the ease or difficulty of prediction. A voting-based scheme was devised to achieve this segregation. Therefore, the pipeline created by Alt et al. [21] based on erudite [76] was followed, and each data sample was scored based on the number of models correctly classifying it. To this end, the input samples were segregated as follows:
@@
The performance analysis highlighted the superiority of the PLM-based models when used in a supervised setting for relation classification. However, under certain conditions, the PLMs’ tendency to memorize the training distribution diffused their focus from the context of the input sample. Thus, this section investigated the context of the text samples to find possible reasons for this behavior. On manually inspecting the samples in the simple and challenging categories, it was found that most samples in the challenging class were difficult to understand, even for a human annotator. Two scenarios were found responsible for this “ambiguity”: 1) when the context provides insufficient or contradictory information, thereby confusing the model, and 2) when the relation definition is too fine-grained, and the model confuses between multiple relation labels. Some examples of the two cases of ambiguity are presented in Table V. The discussion presented below further corroborates these findings.
@@ V-A Ambiguity
The foremost challenge to relation extraction is the aspect of contextual ambiguity in the input data. This issue was apparent from the large percentage of ambiguous samples in the challenging class for supervised relation classifiers. The presence of ambiguous samples with diverging contextual information confuses the models, as the possibility of more than one correct relation exists. Furthermore, it was observed that relations with similar/correlating relation definitions posed a significant challenge for supervised relation extractors. Much like the previous case, these complicated relations led to the possibility of more than one correct relation per sample. Thus, alleviating these ambiguous scenarios would greatly help to make relation extractors more efficient.
On one hand, joint relation extraction presents a natural way of alleviating scenarios arising from ambiguous contexts (detailed discussion in the following section). On the other hand, a probable direction for making relation classifiers robust to this problem could be by employing some input enhancement techniques. Specifically, in cases where the context does not provide sufficient information, data augmentation techniques such as paraphrasing [80] or neural text generation [81] could be employed to expand the context of the text sample and possibly add context and entity-relevant information that can make the relational meaning more apparent. Similarly, for tackling relational ambiguity, more research in the direction of Prompt-PLM methods, which incorporate the verbalizations of the label space in the input prompt, should be conducted. Solving this complex scenario can have many benefits, as it can improve neural networks at fine-grained language understanding.
@@ V-B Joint Relation Extraction
Existing research in this field focuses on testing joint algorithms on a few datasets with similar characteristics. This practice has led to a narrow view of these algorithms, where they perform exceptionally well on the said datasets. According to the findings presented in this paper, one of the significant issues with joint extractors proved to be the false negative problem. This problem stems from the insufficiency of ground-truth labels, so even semantically correct predictions are deemed misclassification. One way to solve this problem would be to create datasets with comprehensive entity and relation annotations. Since this can be a cumbersome task due to the manual efforts involved, using LLMs as relation triplet annotators could be explored as a future direction to create more comprehensive datasets [82]. Second, a more exhaustive solution would be to reformulate the evaluation strategies for joint extractors. A probable avenue could be the introduction of evaluation metrics that do not solely rely on the ground truth labels but evaluate qualitative aspects such as semantic and contextual accuracy of the predicted triplets [83]. Finally, it was found that research on joint relation extraction using LLMs has been scarce in the literature. Thus, conducting exhaustive research in this direction is crucial to making joint relation extractors robust and generalizable.